A 20.0-L volume of methane gas is heated from 25°C to 70°C at constant pressure. What is the final volume of the gas?


1
Expert's answer
2022-05-17T04:52:03-0400

According to Gay Lussac's law, the change in temperature at constant pressure is directly proportional to the volume. that is, V1 / V2 = T1 / T2

we must first transfer the temperature to the kelvin:

T1 = 25+273= 298K

T2 = 70+273= 343K

V1 = 20 L

V1 / V2 = T1 / T2


20/X = 298/343;

X = 20 * 343 / 298 = 23L

ANSWER: 23 L